If 5 people randomly get seated in consecutive seats in a theatre, what is the probability a particular person A sits in the first seat? What is the probability that the first two seats are occupied by A, B in the same order? In some order? What is the probability that A and B sit next to each other?
Answer #1

a) probability a particular person A sits in the first seat =1/5

b)  probability that the first two seats are occupied by A, B in the same order =(1/5)*(1/4) =1/20

c)  probability that A and B sit next to each other =P(considering A and B as one unit and can interchange position within)

=2*4!/(5!)=2/5